# Model-free Legibility: Enhancing Human-Robot Interactions through Implicit Communication and Influence Modulation

### v1

@@ 4.2 Human-Agent Experiment
Figure 2: User study results depicting collaboration and competition performance of humans. The first column shows average success rates, the second column displays violin plots of the final 20 rounds, where the mean values are marked, and the last column visualizes their distributions.
We conducted statistical analysis for the last 20 rounds, including Shapiro-Wilk tests [39] and Bartlett’s tests [40]. The data is normally distributed, but not homoscedastic. Therefore, we used Welch’s t-tests, which is reliable when the samples have unequal variances and possibly unequal sample sizes [41], and Games-Howell Post Hoc tests [42] to verify the significance of the findings. The outcomes from the Games-Howell tests indicate significant disparities in collaboration performance across all human players when engaging with the three distinct types of agents ($p\leq 0.05$). Regarding competition, significant differences were found in human performance when facing Positive-TE agents compared to encounters with Non-TE or Negative-TE agents. However, there was no significant difference ($p\geq 0.05$) in human performance when playing against Non-TE agents compared to interactions with Negative-TE agents.
It is clear that when playing against the Positive-TE agent, human participants achieve greater success rates in both collaboration and competition, when compared to other agents. This seems to be highly desirable behaviour in robots, agents should support humans to achieve their goals in collaboration, and behave transparently (and hence altruistically) when in competition. A Negative-TE agent resists influence, leading to decreased collaboration performance among human players. This results in competitive performance that remains relatively stable when compared to playing against Non-TE agents. Importantly, these real world findings match observations from the simulations conducted in Section 4.1. Figure 3 shows the performances of a trained Non-TE agents when facing the three different types of agents used in the human study. It should be noted that the simulation settings and Q-learning agents are considerably less complex than real human behaviour. Consequently, the effect of the TE reward in these simulations is notably more pronounced than in the user study, though both follow the same trends. Overall, these findings suggest that manipulating influence through TE does indeed impact human behavior. By boosting or resisting influence, we can design agents that facilitate improved or diminished performance in human participants.
@@
Figure 3: Simulation results of Non-TE agents against three agent types. Rows depict collaboration and competition performance. Columns display average success rates, violin plots of the final 20 rounds, where the mean values are marked, and distributions of the last 20 rounds.
Figure 4: Partial survey results of the user study.
@@
After each interaction, participants were asked to give feedback via a short survey about their experience with the agent. Figure 4 presents selected survey findings indicating participants perceive the Pos-TE agent as more legible and human-like. One-way ANOVA tests [43] showed no significant differences suggesting only subtle perceptual distinctions. Combined with the significantly different performance results in Figure 2, this implies that our method enables implicit influence modulation, or implicit communication. For more survey results, refer to Appendix 6.5. Additional findings, limitations and future work is discussed in Appendix 6.7.
5 Conclusion
@@
This work proposed a framework to promote implicit communication for social HRI via influence modulation, which does not require explicit modelling or pre-existing knowledge of human participants. We test the proposed framework in both simulations and human-agent experiments. Results indicate that a positive or negative TE reward can enhance or reduce participant performance based on whether the interaction is collaborative or competitive. Notably, a positive TE reward fosters collaboration and altruism in competition, which is highly desirable for HRI. The framework requires only observations of other participants, and limited domain knowledge or assumptions to produce human-centric behaviours. As a result, this approach has significant potential for broader robotic applications.

### v2

@@ IV-B Human-Agent Experiment
Figure 3: Human-agent experiment results depicting collaboration and competition performance of humans. The first column shows average success rates, the second column displays violin plots of the final 20 rounds, where the mean values are marked, and the last column visualises their distributions.
We conducted a user study [^1] Ethical approval was obtained from Monash University human ethics review board. in a virtual corridor dilemma environment (See Appendix VIII-G for the interface design). Here, human participants played against three agents. Given that humans may display ambivalence, selfishness or altruism, we assumed that a Non-TE agent, being ambivalent, would most accurately represent the typical human. Thus, participants are tested against three agents pre-trained under this assumption: 1. Non-TE agent ($P1$) vs. Non-TE agent ($P2$), 2. Non-TE agent ($P1$) vs. Positive-TE agent ($P2$), and 3. Non-TE agent ($P1$) vs. Negative-TE agent ($P2$). Each participant was asked to play against each of the three pre-trained agents (randomly ordered) for 100 consecutive rounds with randomly generated objectives. The number of rounds was selected to provide ample opportunity for each participant to familiarise themselves with an agent and allow user performance to converge to a stable level. As part of the study, participants were told that one of the opponents could potentially be a real human player. We recruited 26 participants (11 female, 15 male, aged between 20 and 37). Their collaboration and competition performances were observed across 100 rounds of gameplay against each opponent, and subsequently averaged across all participants. These findings are illustrated in Figure 3.
@@
Figure 4: Simulation results of Non-TE agents against three agent types align with the experimental results with humans (Figure 3). Rows depict collaboration and competition performance. Columns display average success rates, violin plots of the final 20 rounds, where the mean values are marked, and distributions of the last 20 rounds.
Figure 5: Selected survey results from the human-agent experiment.
We conducted statistical analysis for the last 20 rounds, including Shapiro-Wilk tests [40] and Bartlett’s tests [41]. The data is normally distributed, but not homoscedastic. Therefore, we used Welch’s ANOVA test [42], which is reliable when the samples have unequal variances and possibly unequal sample sizes, and Games-Howell Post Hoc tests [43] to verify the significance of the findings. The outcomes from the Games-Howell tests indicate significant disparities in collaboration performance across all human players when engaging with the three distinct types of agents ($p\leq 0.05$). Regarding competition, significant differences were found in human performance when facing Positive-TE agents compared to encounters with Non-TE or Negative-TE agents. However, there was no significant difference ($p\geq 0.05$) in human performance when playing against Non-TE agents compared to interactions with Negative-TE agents. The complete statistical report can be found in the Appendix VIII-C.
It is clear that when playing against the Positive-TE agent, human participants achieve greater success rates in both collaboration and competition, when compared to other agents. This seems to be highly desirable behaviour in robots, agents should support humans to achieve their goals in collaboration, and behave transparently (and hence altruistically) when in competition. A Negative-TE agent resists influence, leading to decreased collaboration performance among human players. This results in competitive performance that remains relatively stable when compared to playing against Non-TE agents. Importantly, these real world findings match observations from the simulations conducted in Section IV-A. Figure 4 shows the performances of a trained Non-TE agents when facing the three different types of agents used in the human study. It should be noted that the simulation settings and Q-learning agents are considerably less complex than real human behaviour. Consequently, the effect of the TE reward in these simulations is notably more pronounced than in the user study, though both follow the same trends. Overall, these findings suggest that manipulating influence through TE does indeed impact human behaviour. By boosting or resisting influence, we can design agents that facilitate improved or diminished performance in human participants.
@@
After each interaction, participants were asked to give feedback via a short survey about their experience with the agent. Figure 5 presents selected survey findings indicating participants perceive the Pos-TE agent as more legible and human-like. One-way ANOVA tests [44] showed no significant differences suggesting only subtle perceptual distinctions. Combined with the significantly different performance results in Figure 3, this implies that our method enables implicit influence modulation, or implicit communication. Humans are only moderately aware of differences between policies, despite clear quantitative performance differences. For more survey results, refer to Appendix VIII-H.
IV-C Human-Robot Experiment
The results above show that implicit communication via TE can influence human behaviour in virtual worlds, allowing for greater collaboration and affecting performance in competition, but it is unclear to what extent this holds with a physical robot. To see if the real robot morphology would alter previous results, the corridor dilemma was translated into a physical environment [^2] Ethical approval was obtained from an institutional review board. (Figure 6). We applied the pre-trained policies to a Fetch robot [45]. The robot uses a lidar-based person tracker to track human positions and navigates between discretised grid cells in our pretrained Q table using a timed elastic bands (TEB) planner [46]. More technical details are provided in Appendix VIII-B. The experiment settings remained consistent with those of the human-agent virtual experiment (Section IV-B), with the exception that the grid was not explicitly marked on the floor, and continuous motion used instead of turn-based moves. Participants were instructed to navigate the corridor on foot in a smooth and continuous manner. For each trial, participants began at a randomly assigned starting position and pursued a randomly allocated objective: either to ”meet” (stop in front of the robot) or to ”pass” (avoid ”meet”). Participants were required to walk slowly and steadily towards the finish line, avoiding large strides or lateral movements, while striving to achieve their designated objective. Although the robot navigated using a grid-world-based algorithm, its trajectory was smoothed to generate continuous natural motion.
Figure 6: Human-robot experiment setting. The region measures 2.5 meters in width and extends 6.5 meters in length. Starting positions are labelled as 0 to 4 on the start line. (Refer to the video in the supplementary material for a clearer demonstration.)
Participants interacted with two agents pre-trained under the assumption that a typical human is represented by a Non-TE agent: 1. Non-TE agent ($P1$) vs. Positive-TE agent ($P2$), and 2. Non-TE agent ($P1$) vs. Negative-TE agent ($P2$). We recruited 10 participants (3 women, 7 men, aged between 24 and 34) and asked them to walk towards or past the robot, following randomly generated objectives. Their collaboration and competition performances were observed across 80 rounds of walking using each policy (policies tested successively, order randomised), and subsequently averaged across all participants. These findings are illustrated in Figure 7.
Figure 7: Human-robot experiment results depicting collaboration and competition performance of humans. The first column shows average success rates, the second column displays violin plots of the final 20 rounds, where the mean values are marked.
We performed statistical analysis on the final 20 rounds. The Shapiro-Wilk tests [40] indicated that none of the data followed a normal distribution, while Levene’s test [47] revealed unequal variances in the competition data. Consequently, the Mann-Whitney U test [48] was used to assess the similarity between the ’vs. Pos-TE’ and ’vs. Neg-TE’ data. The results indicate that the SRCL and SRCP distributions for ’vs. Pos-TE’ and ’vs. Neg-TE’ are both significantly different. For a complete statistical report, see Appendix VIII-D.
In general human performance against the physical robot was much better than the results obtained against the virtual agent. Figure 7 shows that participants achieve a substantially higher SRCL when interacting with a Pos-TE robot compared to a Neg-TE robot. This strong effect aligns with our earlier finding that Pos-TE trained robots are desirable for human-robot collaboration.
However, the results also reveal that participants achieve a moderately higher SRCP when interacting with a Neg-TE robot, deviating from findings in previous sections. Several factors could account for this discrepancy. First, we note that the shift away from a more strategic turn-based setting to a continuous motion setting reduces the competitive nature of the game, and introduces a number of potentially confounding variables. Humans are advantaged in the physical world experiment as they are able to move faster, and there are potentially many more variables and factors providing motion cues than the grid-based representation that robot policies used. Moreover, factors such as the physical presence of the robot, and proxemics effects overriding the assigned human objective are also likely to play a role. For example, participants may believe that they have achieved their goal of meeting a robot, while their personal space preferences may lead our system to label the interaction otherwise.
These factors may result in outliers in the relatively small sample sized dataset, in particular for human participants that do not follow our assumption of the average human behaving like a non-TE agent. This is also potentially exacerbated by fatigue, as participants were required to perform 160 trials in an approximately 90 minute session.
Despite these potential confounding factors, it is clear that TE reward augmentation can influence human behaviour and collaboration/ competition in real world interactions with robots. We also evaluated this qualitatively, with participants asked to complete a set of survey questions similar to those in the virtual human-agent experiment. Figure 8 presents selected survey questions. Similar results are obtained to the virtual human-agent experiment survey. Participants perceived the Pos-TE robot as slightly more legible but considerably more human-like. Statistical tests did not reveal significant differences, indicating only subtle perceptual differences with robot morphology. For additional survey results, please refer to Appendix VIII-H.
Figure 8: Partial survey results of the human-robot experiment.
V Discussion
The experiments above validate that the proposed framework can influence human behaviour and increase or decrease human performance. Choosing appropriate TE augmentations to create a situation of information asymmetry for a human-robot interaction is an interesting philosophical decision, depending on the level of agency we want our robots to achieve. In economics and game theory, information asymmetry is a powerful theory explaining the effects of transactions where incomplete information can tip the balance of power in favour of another agent. The experiments above suggest that we can exploit this, potentially using TE to bias our robots towards an Asimov-like reframing [49]:
  • A robot should cede to a human when their objectives conflict.
  • A robot should seek to fulfil its own objectives, as long as these do not conflict with the law above.
It is important to note that the TE objective allows us to build this behaviour into the robot model while also considering other design choices regarding the impacts of information asymmetry in a given setting.
VI Limitations and future Work
Our influence modulation framework, demonstrated in a two-agent setting, has shown potential for fostering complex cooperative and competitive scenarios. A key benefit of the proposed framework is that it is model-free, a simple reward augmentation can enhance or suppress collaboration/ competition performance in human-robot interactions, without needing to explicitly identify influencing factors or needing explicit human behaviour models. However, this limits the interpretability of the robot behaviour, and potentially raises questions about what type of TE reward augmentation should be used for a desired response, which may be dependent on the level of altruism, selfishness or ambivalence displayed by a human collaborator. Future work investigating whether the proposed framework can be extended to infer the type of agent a human most closely resembles (neg-TE, non-TE, pos-TE) may help with appropriate reward augmentation selection.
As is the case with all human-robot experiments, there may be a number of confounding factors influencing our experimental results, from sample size to fatigue. This work lays the groundwork for expanding transfer entropy reward augmentation, but there is still room for significant exploration of these factors in alternative applications and extended settings. We also note that the current Q-learning-based implementation may be limited in scalability to higher dimensional continuous observations and action spaces. To address these limitations, we aim to extend the framework using Gaussian emission neural networks and deep reinforcement learning models, which can handle continuous state and action spaces more effectively. For example, in [25], continuous signals (e.g., velocities, accelerations, positions) were used within a range of probabilistic models to determine distributions over continuous variables and compute transfer entropy. Based on this prior work, we anticipate that the proposed framework will be applicable to continuous signals and enable even more interesting interactions between teams. As suggested by our human-robot experiments, deeper investigations regarding how robot morphology affects influence modulation will be an interesting avenue of future work.
VII Conclusion
@@
This work proposed a framework to promote implicit communication for social HRI via influence modulation, which does not require explicit modelling or pre-existing knowledge of human participants. We test the proposed framework in simulations, human-agent and human-robot experiments. Results indicate that a positive or negative TE reward can enhance or reduce participant performance based on whether the interaction is collaborative or competitive. Notably, a positive TE reward fosters collaboration, which is highly desirable for HRI. Overall, the framework requires only observations of other participants, and limited domain knowledge or assumptions to produce human-centric behaviours. As a result, we believe that this approach has significant potential for broader robotic applications.